OpenStudy (anonymous):

Find a cubic function with the given zeros. 7, -3, 2

OpenStudy (anonymous):

i keep getting x^3+6x^2-13x-42

OpenStudy (anonymous):

but the answer choices r f(x) = x3 - 6x2 - 13x - 42 f(x) = x3 - 6x2 + 13x + 42 f(x) = x3 - 6x2 - 13x + 42 f(x) = x3 + 6x2 - 13 + 42

OpenStudy (anonymous):

x^3+6x^2-13x-42 solves to -7, 3 and -2

OpenStudy (anonymous):

Instead of 7, -3, 2

OpenStudy (anonymous):

f(x) = (x-7)(x+3)(x-2) you have to turn the zeros into factors: x=7 => x-7 = 0 x-7 is a factor

OpenStudy (anonymous):

just have to multiply out to get the cubic you desire

OpenStudy (anonymous):

ohh so i would really be (x-7) (x+3) (x-2)

OpenStudy (anonymous):

im still not getting the answer

OpenStudy (anonymous):

\((x-7)(x+3) =x^2-4x-21\) then multiply by (\(x-2)\)

OpenStudy (anonymous):

so would is get f(x) = x3 - 6x2 + 13x + 42

OpenStudy (anonymous):

nope, check your multiplication

OpenStudy (anonymous):

sorry i was wrong f(x) = x3 - 6x2 - 13x + 42

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

you're welcome